OpenStudy (anonymous):

Finding the derivative of f(x) the long way: f(x)=1/x^2; a=1 formula: lim f(x+h)-f(x)/h as h approaches 0.

OpenStudy (anonymous):

I'm having difficulties solving it the long way. This is the problem \[f(x)=\frac{ 1 }{ x^2 };a=1\] Limit derivative Formula: \[\lim_{h \rightarrow 0}=\frac{ f(x+h)-f(x) }{ h }\]

OpenStudy (anonymous):

I know the answer should be -2/x^3 using the shortcut method, but i'm having difficulty getting to that answer. Here's what I have so far: \[f(x+h)=\frac{ 1 }{ (x+h )^2}\] \[f(x) = 1/x^2\] values plugged into limit formula: \[\lim_{h \rightarrow 0}\frac{ \frac{ 1 }{ (x+h)^2 }-\frac{ 1 }{ x^2 } }{ h }\]

OpenStudy (turingtest):

write the top as a single fraction, simply and expand

OpenStudy (anonymous):

so what i had earlier becomes \[=\lim_{h \rightarrow 0}\frac{ \frac{ 1 }{ x^2+2xh+h^2 }- \frac{ 1 }{ x^2 }}{h }\]

OpenStudy (turingtest):

don't expand until after getting the numerator all as one fraction

OpenStudy (turingtest):

at least i wouldn't recommend it

OpenStudy (turingtest):

i guess it doesn't matter...

OpenStudy (anonymous):

@TuringTest So following your example I get this so far: \[\lim_{h \rightarrow 0}\frac{ \frac{ 1 }{ (x+h)^2 }-\frac{ 1 }{ x^2 } }{h }=\frac{ \frac{ x^2 -(x+h)^2}{ x^2(x+h)^2 } }{ h }\]

OpenStudy (turingtest):

yes, now simplify the denominator and expand the numerator

OpenStudy (turingtest):

by 'simplify the denominator' i mean get that triple fraction into a normal one by putting that bottom h in the denom of the upper fraction

OpenStudy (anonymous):

Don't you mean in the numerator of the upper fraction, aren't I multiplying the entire upper fraction by 1/h?

OpenStudy (turingtest):

i think we mean the same thing, yeah

OpenStudy (turingtest):

\[\lim_{h \rightarrow 0}\frac{ \frac{ 1 }{ (x+h)^2 }-\frac{ 1 }{ x^2 } }{h }=\lim_{h\to 0}\frac{ \frac{ x^2 -(x+h)^2}{ x^2(x+h)^2 } }{ h }=\lim_{h\to0}\frac{ x^2 -(x+h)^2}{ x^2(x+h)^2 } \cdot\frac1h\]

OpenStudy (anonymous):

ok good thats what I thought.

OpenStudy (anonymous):

The above becomes: \[\lim_{h \rightarrow 0}=\frac{ hx^2-h(x+h)^2 }{ x^2(x+h)^2 }\] right?

OpenStudy (turingtest):

erm.. how did h get into the numerator?

OpenStudy (turingtest):

\[\frac ab\cdot\frac cd={ac\over bd}\]

OpenStudy (anonymous):

oh ok. I see, so: \[\lim_{h \rightarrow 0}=\frac{ x^2-(x+h)^2 }{ hx^2(x+h)^2 }\]

OpenStudy (turingtest):

yes, now expand the numerator, simplify, and take the limit

OpenStudy (anonymous):

\[\lim_{h \rightarrow 0}\frac{ x^2-x^2+2xh+h2 }{ hx^2(x+h)^2 }=\frac{ h(2x+h) }{hx^2(x^2+2xh+h^2)} =\frac{ h(2x+h) }{ hx^4+2x^3h^2+h^3x^2 }\] \[=\frac{ h(2x+h) }{ hx(x^3+ 2x^2h+h^2x)} \] this is what I have so far

OpenStudy (turingtest):

wonderful, now cancel what can cancel

OpenStudy (turingtest):

oh you lost your negative sign, btw

OpenStudy (turingtest):

\[\lim_{h \rightarrow 0}=\frac{ x^2-(x+h)^2 }{ hx^2(x+h)^2 }=\lim_{h \rightarrow 0}\frac{ x^2-x^2-2xh-h^2 }{ hx^2(x+h)^2 }\]

OpenStudy (anonymous):

\[=\frac{ -2xh-h^2 }{ hx^2(x^2+2xh+h^2) }=\frac{ -2 }{ hx^2(x^2) }=\frac{ -2 }{ x^4h }\] I think I made a mistake somewhere, cause evaluating at h->0 would give me an undefined answer.